a^3-x * b^5x=a^x+5 * b^3x,find the value of x log(b/a) Ans log a
Please give the solution

Grade:10

2 Answers

Ajay Verma
10 years ago
solution:


a(3-x) b(5x)= a(x+5)b(3x)
b(5x)/ b(3x) = a(x+5)/ a(3-x)
b(2x)= a(2x+2)
b(2x)= a2x *a2
b(2x)/a2x = a2
take log

log (b(2x)/a2x) = log (a2)
2x log(b/a) = 2 log a

so x log(b/a) = log a


Thanks and Regards,
Ajay verma,
askIITians faculty,
IIT HYDERABAD
Arpit Kumar
10 years ago
Thanks sir